### v2

@@ 5.4 Evaluation
To assess the performance of the proposed approaches, we address the following research questions:
  1. (a)
@@
  2. (b)
  3. (c)
  4. (d)
In Approach 1, we focused on fine-tuning all the transformer models without the use of summarization or augmentation. During training, we incorporated translated fake news (English-Bengali) alongside Bengali news articles. The performance of these fine-tuned models was evaluated using ‘Test Dataset-1’ and ‘Test Dataset-3’. We show the performance of the fine-tuned models on ‘Test Dataset-1’ and ‘Test Dataset-3’ in Table 9 and 11 respectively. The fine-tuned models in Approach 1 exhibited impressive performance on ‘Test Dataset-1’, with the BanglaBERT model achieving the lowest accuracy of 88%. The mBERTBase model achieved the highest accuracy of 92%, while the other three models also scored above 90%. On the other hand, the BanglaBERT Base model attained the highest accuracy score of 80% on ‘Test Dataset-3’. This custom test dataset, designed to assess generalization performance, was completely unfamiliar to the fine-tuned models. The results obtained from Approach 1 indicate that it is feasible to develop a robust Bengali fake news detection model without relying on summarization or augmentation techniques, particularly when a substantial amount of fake data is available. In our case, we achieved this by translating fake English news into Bengali, thereby increasing the number of fake news instances. Overall, the BanglaBERT Base model demonstrated better performance on both test datasets.
@@
Method Model A P R F1 MCC ROC
@@
We introduced the utilization of summarization technique in Approach 2. Analyzing the results presented in Table 9 and 11, we can observe that the BanglaBERT Base model achieved the highest accuracy scores of 88% and 78% on ‘Test Dataset-1’ and ‘Test Dataset-3’ respectively. The DB-mBERT model exhibited good performance on ‘Test Dataset-1’ with an accuracy of 88%, but its performance on ‘Test Dataset-3’ was comparatively weaker, with an accuracy of only 70%. When compared to the results of Approach 1, we observed a slight decrease in performance after implementing the summarization technique in Approach 2. However, this decrease is negligible, with only a marginal 1-2% reduction in accuracy.
@@
Method Model A P R F1 MCC ROC
@@
Method Model A P R F1 MCC ROC
@@ 5.5 Result Analysis
This section seeks to experimentally support the models’ performance. Among the three test datasets, we use the first two test datasets - ‘Test Dataset-1’ and ‘Test Dataset-2’ for quantitative analysis. Table 9 and 10 present the experimental findings that followed the evaluation of each model on these two test datasets. Figure 4 is the visual representation of the performance comparison among all the fine-tuned models across all the approaches on ‘Test Dataset-1’.
@@
(b) Comparison in terms of F1-score.
@@
(b) Comparison in terms of F1-score.
@@
In this section, we evaluate the generalization ability of the model towards real-world test samples using ‘Test Dataset-3’. The results of the experiments on Test Dataset-3 are presented in Table 11. Even on this completely unseen test dataset, Approach 3 outperforms the other three approaches and achieves the highest accuracy and F1 score of 0.86. The equality of the accuracy and F1 score indicates that the model is capable of correctly classifying both positive and negative classes. Figure 6 provides a visual representation of the performance comparison of all models on ‘Test Dataset-3’.
@@
(b) Comparison in terms of F1-score.
@@
Models Accuracy F1-score
@@
Acc F1 Acc F1 Acc F1 Acc F1
@@ 5.6 Comparison with Existing Models
To evaluate and compare the performance of our proposed approaches, we conducted direct inference on two existing pre-trained models: TM-mBERT and DB-mBERT which are specifically designed for detecting Bengali fake news. Table 13 shows result comparison. We use ‘Test Dataset-3’ for performance comparison between existing models and our approaches, because this test dataset is unknown to all the models. All our approaches performed better than the existing pre-trained models. We also conducted experiments using seven traditional machine learning models, including Support Vector Classifier (SVC), K-Nearest Neighbor (KNN), Decision Trees (DT), Random Forest (RF), AdaBoost (ADB), Multinomial Naive Bayes (MNB), and Logistic Regression (LR). Furthermore, we explored three deep learning techniques: Long Short Term Memory network (LSTM), Bidirectional Long Short Term Memory network (Bi-LSTM), and Convolutional Hierarchical Attention Network (Conv HAN). The experimental results are summarized in Table 14. We trained these models using the ‘Dataset 2’ training dataset, maintaining an 85:15 training-to-validation split. Default hyperparameters were employed during training for all models. Textual data were numerically represented using Term Frequency - Inverse Document Frequency (TF-IDF) for traditional models and the Keras Embedding Layer for deep learning models. Testing was conducted using the ‘Test Dataset-3’. According to the findings presented in Table 14, traditional machine learning models, except RF, ADB, and LR, demonstrated their best performance when augmentation was applied before training (Approach 3). RF and ADB attained the highest accuracy scores through summarization before training (Approach 2). Conversely, LR exhibited optimal performance without summarization or augmentation (Approach 1). Among the deep learning models, LSTM achieved the highest accuracy of 81.8% when summarization was employed before training (Approach 2), while Bi-LSTM attained 78.9% accuracy by employing both augmentation and summarization before training (Approach 4). However, akin to RF, Conv-HAN achieved its highest accuracy without any augmentation or summarization (Approach 1). Nonetheless, none of the traditional or deep learning models surpassed the performance of our fine-tuned mBERT-Base on ‘Test Dataset-3’, as mBERT-Base achieved an accuracy of 86% using augmentation (Approach 3).
@@ 6 Conclusion and Future Work
In this paper, we presented our approach to classify Bengali fake news using a combination of summarization and augmentation techniques with pre-trained language models. To ensure a thorough evaluation of our trained models, we conducted tests on three distinct test datasets. The results revealed that our models achieved remarkably high levels of accuracy and f1-score on the first two test datasets. For the third test dataset, which was kept separate to assess generalization, the best model demonstrated an accuracy and f1-score of 86%. These findings demonstrate the effectiveness of the model in accurately distinguishing between fake and authentic news articles. There are several promising directions for future work that can expand upon our research and broaden the scope of this paper. While our primary focus was on the classification of Bengali fake news, our approach holds potential for application in other languages with limited resources. Furthermore, our research primarily centered around the binary classification of authentic and fake news articles. To further advance the field, future studies could explore the realm of multi-class classification, involving the categorization of news articles into different types of fake news, including satire, propaganda, and clickbait.
@@ Appendix A CustomFake Corpus Creation
(d) Annotation Quality. In response to the first question, the F1-scores for determining the fake class among the three annotators are as follows: 65%, 70%, and 63% respectively. The inter-annotator agreement, assessed using Fleiss’ Kappa fleiss1971measuring stands at 44.23%, indicating that our human annotators provided identical responses in nearly 44% of the cases. Regarding the second question, when the news is authentic, an average of 72% of respondents select ‘The content is believable’, while 20% opt for ‘Source is reliable’. Here, the term ‘source’ refers to an individual or organization capable of validating the reported news. Conversely, when the news is deemed fake, an average of 51% of respondents indicate ‘The content is unrealistic’, while 39% suggest ‘Has no trustworthy source’ as the reason.

### v3

@@ 5.4 Evaluation
The research questions outlined in Section 1 are discussed in detail below.
@@
In Approach 1, we focused on fine-tuning all the transformer models without the use of summarization or augmentation. During training, we incorporated translated fake news (English-Bengali) alongside Bengali news articles. The performance of these fine-tuned models was evaluated using ‘Test Dataset-1’ and ‘Test Dataset-3’. We show the performance of the fine-tuned models on ‘Test Dataset-1’ and ‘Test Dataset-3’ in Table 8 and 10 respectively. The fine-tuned models in Approach 1 exhibited impressive performance on ‘Test Dataset-1’, with the BanglaBERT model achieving the lowest accuracy of 88%. The mBERTBase model achieved the highest accuracy of 92%, while the other three models also scored above 90%. On the other hand, the BanglaBERT Base model attained the highest accuracy score of 80% on ‘Test Dataset-3’. This custom test dataset, designed to assess generalization performance, was completely unfamiliar to the fine-tuned models. The results obtained from Approach 1 indicate that it is feasible to develop a robust Bengali fake news detection model without relying on summarization or augmentation techniques, particularly when a substantial amount of fake data is available. In our case, we achieved this by translating fake English news into Bengali, thereby increasing the number of fake news instances. Overall, the BanglaBERT Base model demonstrated better performance on both test datasets.
@@
Method Model A P R $\ignorespaces\mathbf{F}_{1}$ MCC ROC
@@
We introduced the utilization of summarization technique in Approach 2. Analyzing the results presented in Table 8 and 10, we can observe that the BanglaBERT Base model achieved the highest accuracy scores of 88% and 78% on ‘Test Dataset-1’ and ‘Test Dataset-3’ respectively. The DB-mBERT model exhibited good performance on ‘Test Dataset-1’ with an accuracy of 88%, but its performance on ‘Test Dataset-3’ was comparatively weaker, with an accuracy of only 70%. When compared to the results of Approach 1, we observed a slight decrease in performance after implementing the summarization technique in Approach 2. However, this decrease is negligible, with only a marginal 1-2% reduction in accuracy.
@@
Method Model A P R $\ignorespaces\mathbf{F}_{1}$ MCC ROC
@@
Method Model A P R $\ignorespaces\mathbf{F}_{1}$ MCC ROC
@@ 5.5 Result Analysis
This section seeks to experimentally support the models’ performance. Among the three test datasets, we use the first two test datasets - ‘Test Dataset-1’ and ‘Test Dataset-2’ for quantitative analysis. Table 8 and 9 present the experimental findings that followed the evaluation of each model on these two test datasets. Figure 6 is the visual representation of the performance comparison among all the fine-tuned models across all the approaches on ‘Test Dataset-1’.
@@
(b) Comparison in terms of $\ignorespaces F_{1}$ -score.
@@
(b) Comparison in terms of $\ignorespaces F_{1}$ -score.
@@
In this section, we evaluate the generalization ability of the model towards real-world test samples using ‘Test Dataset-3’. The results of the experiments on Test Dataset-3 are presented in Table 10. Even on this completely unseen test dataset, Approach 3 outperforms the other three approaches and achieves the highest accuracy and $\ignorespaces F_{1}$ score of 0.86. The equality of the accuracy and $\ignorespaces F_{1}$ score indicates that the model is capable of correctly classifying both positive and negative classes. Figure 8 provides a visual representation of the performance comparison of all models on ‘Test Dataset-3’.
@@
(b) Comparison in terms of $\ignorespaces F_{1}$ -score.
@@
Models Accuracy $\ignorespaces\mathbf{F}_{1}$ -score
@@
Acc $\ignorespaces\mathbf{F}_{1}$ Acc $\ignorespaces\mathbf{F}_{1}$ Acc $\ignorespaces\mathbf{F}_{1}$ Acc $\ignorespaces\mathbf{F}_{1}$
@@
5.6 Misclassification Analysis
We conducted a brief error analysis on our best-performing model (mBERT Base - Approach 3) using Test Dataset-3. Out of 204 instances, it misclassified one real news article and 28 fake news articles. Table 14 shows examples of misclassified instances.
(a) False Negative: In the single false negative case, the model misclassified a real news article as fake, assigning it a very low confidence score (0.0047), indicating high uncertainty in its prediction. Our analysis revealed the following contributing factors:
  1. 1.
    Repetition of certain word: The repeated use of the word “guNJ/jn[rumor] may have misled the model into associating the article with speculation and misinformation.
  2. 2.
    Political Context: The article discussed upcoming elections and political figures, which may have triggered features commonly associated with fabricated political narratives.
  3. 3.
    Entertainment Context: Since the article involved a popular actor, the presence of entertainment-related content could have contributed to the model’s confusion, as entertainment articles are often exaggerated or fictionalized.
(b) False Positive: Out of 28 false positive cases—fake news instances that were incorrectly predicted as real—we identified several common characteristics:
  1. 1.
    Numerical evidences: Articles containing precise statistics, monetary values, or measurements (e.g., "21 crore USD") often appeared more credible to the model. This reliance on quantitative data as a proxy for truth led to misclassification.
  2. 2.
    Long texts: Several misclassified fake articles were lengthy and detailed, which possibly increased their perceived credibility. The model may have associated longer texts with more thorough reporting, thus mistakenly classifying them as real.
  3. 3.
    Other sources as evidence: Some fake articles attempted to build authenticity by mentioning other sources, organizations, or timelines. This inclusion of evidence-like elements misled the model, as it lacked the ability to verify these claims.
  4. 4.
    Lack of external reasoning: BERT-based models are inherently limited in understanding real-world plausibility. As a result, even highly implausible or factually inaccurate articles were misclassified if their surface features aligned with real news patterns. For instance, stories about supernatural events or implausible medical conditions (e.g., a child growing animal fur) were predicted as real simply due to formal language and structured storytelling.
News (Translated) True Label Predicted Label
’pRdhanmn/tRii caIel AbshYI inr/bacn krb. baNNGlaedsh clic/cetRr jnipRy nayk ephredousek Ebar edkha eJet paer jatiiy sNNGsd inr/bacen, Emn guNJ/jn eshana Jaic/chl imiDya parhay. sm/pRit itin pRdhanmn/tRii eshkh Haisnar seNG/g raSh/TRiiy spher ghuer Eesechn. Aar Erpr ethekI EI guNJ/jenr shuru. Ebar EI guNJ/jenr jbab ideln ephredous. edesh ipherI inr/bacn pRseNG/g EkiT sNNGbad madhYem mukh khuleln EI Aibhenta. ephredous beln, Ekhena ikchu phaInal Hyin. Emn ktha AaimO shunich. sbarI menabasna thaek. AamarO Aaech. jaygaTa EkhnO OIrkm Abs/thaetI Aaech. EkhnO ikchuI Hyin. Aamaek Jid na Aphar ker Aaim HNNYa ba na bleba ekmn ker. Aamaek EkhnO dl ethek Aphar kra Hyin. Aphar na krel ejar ker ikchu kra Jaeb na. dl ethek pRs/tab edOya Hel AbshYI inr/bacn krb. itin AarO beln, Aaim ik bleba buejh UThet parich na. Aaim pRdhanmn/tRiir seNG/g spher igeyich. ETa Aamar Aenk brh paOya. Erpr Uin Jid Aamaek iney ikchu ebheb thaekn taHel AalHamduillLaH. Aaim isenmar manuSh. kaejr madhYem manuShek ibenadn ideyich. Ekhn Aamar dWara Jid edsh O deshr Upkar Hy taHel AbshYI Aaim edsh O deshr jnY kaj kret caI. manueShr esbay Jid inejek ineyaijt kret pair Er ecey esoubhaegYr Aar ikchu enI. [E] "If the Prime Minister wants, I will definitely run in the election." Rumors were circulating in the media that popular Bangladeshi film actor Ferdous might contest in the upcoming national elections. Recently, he accompanied Prime Minister Sheikh Hasina on a state visit, which fueled the speculation. Upon returning, Ferdous addressed the rumors in an interview, saying, "Nothing has been finalized yet. I’ve heard the rumors too. Everyone has their aspirations, and so do I. The situation is still in flux. If I am not offered a candidacy, how can I say yes or no? I haven’t been offered anything by the party yet. If they do, I will definitely run." He added, "I don’t know what to say. Traveling with the Prime Minister was a great honor. If she sees potential in me, Alhamdulillah. I am from the film industry and have entertained people through my work. If I can serve the nation, there’s no greater honor." 1 0
cNNapaInbabgNJ/j ejlay tasiphJa namk saerh itn bchr bJsii ishshur gaey pshur elam UThech. cNNapaInbabgeNJ/j ibrl eraeg AakRan/t dirdR pirbaer saerh itn bcherr ishshu knYa tasiphJa jaHan muinra. tar baba idnmjur masuduj/jaman mamun. bairh naecal bnibbhaegr paesh egaDaUn paDa.ishshu tasiphJa jenMr pr ethekI tar shriierr lmWa lmWa pshm edkha Jay. idn JtI grhaec/ch pshmguilO barhet barhet pshur meta edkha Jaec/ch. ishshu tasiphJar shriierr ipeThr echaT/T EkiT iTUmar ethek EiTr Ut//pit/t bel tasiphJar ma tanijla khatun janan. [E] In Chapainawabganj district, a three-and-a-half-year-old girl named Tasfia is growing "animal-like" hair on her body. She suffers from a rare condition. Her father, Masuduzzaman Mamun, is a day laborer, and they live in Godown Para, near the Nachol Forest Department. Since birth, long hair has been growing on Tasfia’s body. As time passes, the hair is growing thicker, giving her an animal-like appearance. According to her mother, Tanzila Khatun, the condition appears to originate from a small tumor on her back. 0 1
Table 14: Misclassification analysis example. Label 0 denotes fake and 1 denotes real.
@@ 5.7 Comparison with Existing Models
To evaluate and compare the performance of our proposed approaches, we conducted direct inference on two existing pre-trained models: TM-mBERT and DB-mBERT which are specifically designed for detecting Bengali fake news. Table 12 shows result comparison. We use ‘Test Dataset-3’ for performance comparison between existing models and our approaches, because this test dataset is unknown to all the models. All our approaches performed better than the existing pre-trained models.
We also conducted experiments using seven traditional machine learning models, including Support Vector Classifier (SVC), KNN, DT, RF, AB, MNB, and LR. Furthermore, we explored three deep learning techniques: Long Short Term Memory network (LSTM), Bidirectional Long Short Term Memory network (Bi-LSTM), and Convolutional Hierarchical Attention Network (Conv HAN). We trained these models using the ‘Dataset 2’ training dataset, maintaining a 90:10 training-to-validation split and testing was conducted using the ‘Test Dataset-3’. Default hyperparameters were employed during training for all models. The experimental results are summarized in Table 13.
For traditional machine learning models, textual data were numerically represented using Term Frequency-Inverse Document Frequency (TF-IDF). For deep learning approaches, we first preprocessed text using TensorFlow Keras’ Tokenizer with a 10,000-word vocabulary, converting sequences to fixed-length 100-token representations through post-padding. All neural models (LSTM, BiLSTM, and Conv-HAN) employed 128-dimensional trainable embeddings. The Conv-HAN architecture specifically integrated 1D convolutions (64 filters, kernel size=3), max-pooling, an LSTM layer, and hierarchical attention mechanisms. Each deep learning model utilized a sigmoid output layer and was trained for 5 epochs (batch size=32) using Adam optimization, with consistent hyperparameters across architectures.
According to the findings presented in Table 13, traditional machine learning models, except RF, ADB, and LR, demonstrated their best performance when augmentation was applied before training (Approach 3). RF and ADB attained the highest accuracy scores through summarization before training (Approach 2). Conversely, LR exhibited optimal performance without summarization or augmentation (Approach 1). Among the deep learning models, LSTM achieved the highest accuracy of 81.8% when summarization was employed before training (Approach 2), while Bi-LSTM attained 78.9% accuracy by employing both augmentation and summarization before training (Approach 4). However, akin to RF, Conv-HAN achieved its highest accuracy without any augmentation or summarization (Approach 1). Nonetheless, none of the traditional or deep learning models surpassed the performance of our fine-tuned mBERT-Base on ‘Test Dataset-3’, as mBERT-Base achieved an accuracy of 86% using augmentation (Approach 3).
@@ 6 Conclusion and Future Work
In this paper, we presented our approach to classify Bengali fake news using a combination of summarization and augmentation techniques with pre-trained language models. To ensure a thorough evaluation of our trained models, we conducted tests on three distinct test datasets. The results revealed that our models achieved remarkably high levels of accuracy and $\ignorespaces F_{1}$ -score on the first two test datasets. For the third test dataset, which was kept separate to assess generalization, the best model demonstrated an accuracy and $\ignorespaces F_{1}$ -score of 86%. These findings demonstrate the effectiveness of the model in accurately distinguishing between fake and authentic news articles. There are several promising directions for future work that can expand upon our research and broaden the scope of this paper. While our primary focus was on the classification of Bengali fake news, our approach holds potential for application in other languages with limited resources. Furthermore, our research primarily centered around the binary classification of authentic and fake news articles. To further advance the field, future studies could explore the realm of multi-class classification, involving the categorization of news articles into different types of fake news, including satire, propaganda, and clickbait. Additionally, future work could explore the use of generative models for data augmentation, such as BigBird, which can handle longer documents and capture document-level semantics. Moreover, evaluating newer large language models (LLMs), which continue to evolve rapidly in terms of performance and adaptability, could lead to more robust and context-aware fake news classification systems.
@@ Appendix A CustomFake Corpus Creation
(d) Annotation Quality. In response to the first question, the $\ignorespaces F_{1}$ -scores for determining the fake class among the three annotators are as follows: 65%, 70%, and 63% respectively. The inter-annotator agreement, assessed using Fleiss’ Kappa fleiss1971measuring stands at 44.23%, indicating that our human annotators provided identical responses in nearly 44% of the cases. Regarding the second question, when the news is authentic, an average of 72% of respondents select ‘The content is believable’, while 20% opt for ‘Source is reliable’. Here, the term ‘source’ refers to an individual or organization capable of validating the reported news. Conversely, when the news is deemed fake, an average of 51% of respondents indicate ‘The content is unrealistic’, while 39% suggest ‘Has no trustworthy source’ as the reason.